### v2

# UnlearnCanvas: A Stylized Image Dataset to Benchmark Machine Unlearning for Diffusion Models

Abstract
The rapid advancement of diffusion models (DMs) has not only transformed various real-world industries but has also introduced negative societal concerns, including the generation of harmful content, copyright disputes, and the rise of stereotypes and biases. To mitigate these issues, machine unlearning (MU) has emerged as a potential solution, demonstrating its ability to remove undesired generative capabilities of DMs in various applications. However, by examining existing MU evaluation methods, we uncover several key challenges that can result in incomplete, inaccurate, or biased evaluations for MU in DMs. To address them, we enhance the evaluation metrics for MU, including the introduction of an often-overlooked retainability measurement for DMs post-unlearning. Additionally, we introduce UnlearnCanvas, a comprehensive high-resolution stylized image dataset that facilitates us to evaluate the unlearning of artistic painting styles in conjunction with associated image objects. We show that this dataset plays a pivotal role in establishing a standardized and automated evaluation framework for MU techniques on DMs, featuring 7 quantitative metrics to address various aspects of unlearning effectiveness. Through extensive experiments, we benchmark 5 state-of-the-art MU methods, revealing novel insights into their pros and cons, and the underlying unlearning mechanisms. Furthermore, we demonstrate the potential of UnlearnCanvas to benchmark other generative modeling tasks, such as style transfer. The UnlearnCanvas dataset, benchmark, and the codes to reproduce all the results in this work can be found at https://github.com/OPTML-Group/UnlearnCanvas.
Machine Learning
1 Introduction
In recent years, the field of text-to-image technology, epitomized by Diffusion Models (DMs), has seen remarkable progress, demonstrating an unprecedented ability to generate high-resolution, high-quality images across a wide array of subjects (Rombach et al., 2022; Midjourney et al.,, 2023). These models, widely accessible to public, find applications in diverse sectors, including advertising, the creative arts (Fotor et al.,, 2022), and forensic sketching (LabLab AI et al.,, 2023). Central to the training of DMs is their reliance on the data source from the Internet (Schuhmann et al., 2022) to cover a wide range of content. However, this inclusiveness comes with challenges, including potential concerns about harmful generation (Schramowski et al., 2023), copyright issues (Krietzberg, 2023), and biases/stereotypes (Vinker et al., 2023; Turk, 2023).
Figure 1: An illustration of the task MU for DMs. The pretrained model contains the generation capability of different concepts in different domains. MU aims to erase the generation of a certain unlearning target concept without harming the generation of others.
To alleviate the negative social impacts associated with DMs, machine unlearning (MU) techniques are catching increasing attention and have been studied in the field of vision generative models (Schramowski et al., 2023; Gandikota et al., 2023a, b; Kumari et al., 2023; Fan et al., 2023; Heng & Soh, 2023). MU for DMs (see Fig. 1 for an illustration) aims to erase the model’s capacity to generate images related to a specific concept, essentially erasing that concept from the pre-trained model. This concept can be provided in the form of text prompts (i.e., text conditions of DMs), covering a wide range from concrete object entities to abstract notions, such as artistic styles. We refer readers to Sec. 2 for more related work on MU for DMs.
Unlike MU studies on image classification (Izzo et al., 2021; Golatkar et al., 2020; Becker & Liebig, 2022; Thudi et al., 2022; Warnecke et al., 2021; Jia et al., 2023), we lack evaluation benchmarks for MU in DMs. The current primary choice for assessing the performance of unlearned DMs (i.e., DMs post unlearning) is the Inappropriate Image Prompts (I2P) dataset (Schramowski et al., 2023). This dataset consists of real-world image-to-text prompts designed for erasing concepts like ‘nudity’ and ‘violence’. However, I2P mainly contains isolated unlearning targets and focuses on safety issues related to image generation.
To enhance the assessment of MU in DMs and establish a standardized evaluation framework, we propose to develop a new benchmark dataset, referred to as UnlearnCanvas, which is designed to evaluate the unlearning of artistic painting styles along with associated image objects. Unlike I2P, UnlearnCanvas can be used to assess the removal of copyrighted information, such as artistic style, and it allows for increased diversity by including various object classes and styles. We summarize our contributions below.
$\bullet$ We conduct a systematic review of existing MU methods for DMs and identified three unresolved challenges in their evaluations, such as limited quantitative assessments and a lack of retainability evaluation of unlearned DMs.
$\bullet$ We propose UnlearnCanvas, a large-scale high-resolution stylized image dataset with 24,000 labeled images, 60+ styles, and 20+ object topics. Its style-object dual supervision and high stylistic consistency help us to improve the precision of MU evaluation for DMs.
$\bullet$ We develop a standardized evaluation pipeline for MU in DMs under UnlearnCanvas, which measures unlearning effectiveness, generation retainability, generation quality, and unlearning efficiency.
$\bullet$ We benchmark state-of-the-art MU methods for DMs using UnlearnCanvas, validating the significance of our proposed evaluation metrics and providing insights into the strengths, weaknesses, and unlearning mechanisms.
2 Related Work and Preliminaries
Machine unlearning (MU) for generative models in vision. MU was originally developed to mitigate the (potentially detrimental) influence of specific data points in a pretrained ML model, without necessitating a complete retraining of the model after removing these unlearning data (Cao & Yang, 2015; Bourtoule et al., 2021; Xu et al., 2023). The significance of MU has emerged with the purpose of data privacy protection, in response to regulations such as ‘the right to be forgotten’ (Hoofnagle et al., 2019). However, it has rapidly gained recognition for its role in promoting security, safety, and trustworthiness of ML models. Examples include defense against backdoor attacks (Liu et al., 2022; Jia et al., 2023), fairness enhancement (Chen et al., 2023; Oesterling et al., 2023), and controllable federated learning (Halimi et al., 2022; Jin et al., 2023).
More recently, unlearning has proven crucial in text-to-image generation via DMs to prevent the production of harmful, private, or illegal content (Gandikota et al., 2023a; Zhang et al., 2023a; Gandikota et al., 2023b; Fan et al., 2023; Kumari et al., 2023; Heng & Soh, 2023). For example, DMs have demonstrated a susceptibility to generating NSFW (Not Safe For Work) images when conditioned on inappropriate text prompts (e.g., ‘nudity’ and ‘violence’) (Schramowski et al., 2023; Zhang et al., 2023b; Chin et al., 2023). Similarly, MU has also found applications in preventing the generation of copyrighted information, such as the misuse of artistic painting styles (Gandikota et al., 2023a; Zhang et al., 2023a). In summary, MU can be regarded as a (post-training) model edit operation, aimed at preventing DMs from producing undesired or inappropriate images.
Our setup and objective. Let us denote a pretrained DM by $\bm{\theta}_{\mathrm{o}}$, capable of synthesizing high-quality images conditioned on a provided text prompt $c$ (e.g., ‘A painting of a cat in Van Gogh style’). In this work, we focus on the task of generating images in various painting styles. When there is a request to adapt the model $\bm{\theta}_{\mathrm{o}}$ to prevent the generation of images, e.g., in the ‘Van Gogh’ style, the MU problem arises, with the ‘Van Gogh’ style being the specified unlearning target, also known as the erasing concept (Gandikota et al., 2023a). See Fig. 1 for a visual representation of the concept.
In formulating MU methods, the traditional setup presupposes access to the training dataset $\mathcal{D}$, where $\mathcal{D}_{\mathrm{f}}\subset\mathcal{D}$ represents the forget set that revolves around the unlearning target. A commonly used gold standard of MU is to retrain the model from scratch over the retain set $\mathcal{D}_{\mathrm{r}}=\mathcal{D}\setminus\mathcal{D}_{\mathrm{f}}$. However, commencing the training process from scratch becomes practically unfeasible because of (1) the substantial computational expense it entails, (2) the difficulty in accurately identifying $\mathcal{D}_{\mathrm{f}}$, and (3) the absence of access to the original training data source.
To overcome the aforementioned difficulties (1)-(3), the current research mainstream (Gandikota et al., 2023a; Kumari et al., 2023; Zhang et al., 2023a; Gandikota et al., 2023b; Heng & Soh, 2023; Fan et al., 2023) has focused on developing approximate yet efficient unlearning methods as practical alternatives to full retraining. The approximation introduced by current unlearning methods raises the question that we aim to address in this work:
How can we accurately assess unlearning performance and establish a benchmark for existing unlearning methods?
3 Unresolved Challenges in MU Evaluation
This work will develop a new benchmark dataset, referred to as UnlearnCanvas (detailed in Sec. 4), focusing on the task of unlearning artistic painting styles in conjunction with associated image objects. With the use of UnlearnCanvas, a text prompt for DMs can be constructed as ‘A painting of a {object} in {style}’. And the unlearning target can be explicitly designated as either ‘style’ or ‘object’. To be specific, UnlearnCanvas is targeted for the open challenges (C1-C3) when building a rigorous quantitative evaluation framework for MU in DMs. These challenges are consolidated through a careful examination of existing studies for object and/or style unlearning in Tab. 1.
Table 1: An overview of the scope and evaluation methods of the existing MU methods for DMs. The symbol denotes neither objects nor styles are covered.
Related work Unlearning targets Quantitative evaluation
Unlearning
effectiveness
Retainability
ESD (Gandikota et al., 2023a)
5 artistic styles
11 objects
objects only objects only
CA (Kumari et al., 2023)
4 artistic styles
4 objects
objects only
FMN (Zhang et al., 2023a)
1 artistic style
7 objects
objects only
UCE (Gandikota et al., 2023b)
7 artistic settings [^1] UCE adopted a multi-concept unlearning setting, where different numbers of styles are unlearned at the same time and total up to 1000 styles are considered.
11 objects
objects only objects only
SA (Heng & Soh, 2023)
1 artistic style
31 objects
objects only
SalUn (Fan et al., 2023)
No artistic styles
11 objects
objects only objects only
(C1) The absence of a diverse yet cohesive unlearning target repository. As shown in Tab. 1, the assessment of MU for DMs, in terms of both effectiveness (concept erasure performance) and retainability (preserved generation quality under non-forgotten concepts), is typically conducted using manually selected targets, often chosen from a limited pool of unlearning targets. A critical aspect missing from MU evaluation is the recognition that unlearning targets related to styles and objects should not be isolated within the dataset used for MU assessment. Instead, they should be considered in a manner that exposes their potential interconnected influences on other targets.
Table 2: Classification accuracy of a style classifier (ViT-L/16) finetuned on WikiArt (Phillips & Mackintosh, 2011). The classifier is tested on the test set and the image set generated by SD v1.5 and v2.0 with the prompt ‘An painting in {artist} style’, where the {artist} is traversed over the WikiArt labels (10 images/prompt).
WikiArt
Train Set
WikiArt
Test Set
Images by
SD v1.5
Images by
SD v2.0
Accuracy 99.8% 85.4% 56.7% 41.2%
(C2) The lack of precision in evaluation. Artistic styles can be challenging to precisely define and distinguish, which makes quantifying their impact on unlearning effectiveness and retainability difficult. To illustrate this challenge, we conducted a preliminary investigation in Tab. 2 using a ViT-based classifier trained on the WikiArt dataset (Phillips & Mackintosh, 2011). This dataset categorizes paintings into 139 distinct styles, each annotated with the creator’s name. Despite training, the ViT-large model achieved only an $84.32\%$ accuracy on the test dataset. This result underscores the inherent complexity of classifying artistic styles, causing inaccuracies in quantitative MU evaluations. The classifier’s suboptimal accuracy on DMs generated images further increases the prediction error (around $50\%$ accuracy). This complicates precise quantitative evaluations.
(C3) The lack of a systematic study on ‘retainability’ of DMs post unlearning. As indicated by the ‘retainability’ column in Tab. 1, assessing the capacity of unlearned DMs to maintain image generation under innocent concepts is notably lacking. However, retainability is a critical indicator of potential ‘side effects’ of an unlearning method. In conjunction with challenge (C1), the inclusion of diverse unlearning targets, encompassing both styles and objects, allows for the evaluation of retainability from various perspectives. For instance, when unlearning an artistic style, it is crucial to consider the retainability of image generation abilities for other styles as well as for objects.
The challenges (C1)-(C3) outlined above have driven us to develop UnlearnCanvas as a solution for the systematic and comprehensive evaluation of MU in DMs. We will show in Sec. 4 that UnlearnCanvas includes a diverse range of artistic styles and objects. Within this dataset, we can explore the interrelationships between paired image styles and image objects to address (C1). Each style category in UnlearnCanvas maintains high intra-style consistency, facilitating accurate style and object classification and enabling post-unlearning quantitative evaluation for (C2). To address (C3), we can evaluate two types of retainability, providing a precise and comprehensive assessment of the potential side effects induced by MU methods.
4 Our Proposal: UnlearnCanvas Dataset
Construction of UnlearnCanvas. As motivated in Sec. 2, UnlearnCanvas is created for ease of MU evaluation in DMs. Its construction involves two main steps: seed image collection and subsequent image stylization; see Fig. 2 for a schematic overview.
Seed image collection. UnlearnCanvas includes images in $60$ unique artistic styles across $20$ distinct objects. The style categories are derived from a set of high-resolution photos, termed seed images. These seed images consist of original real-world photos from Pexels (Pexels et al.,, 2023). In total, there are 400 seed images across 20 object classes.
Image stylization. After collecting the seed images, the next step in dataset development is the stylization process. During this phase, each seed image is transformed into 60 predetermined artistic styles. The goal is to ensure that the resulting images maintain high stylistic consistency while preserving the content details present in the seed images. This stylization process is carried out using the services provided by (Fotor et al.,, 2022). Once all the seed images have been stylized in all 60 styles, the dataset is constructed with super-high-resolution images and organized in a hierarchical structure that balances both style and object categories.
Figure 2: An illustration of the key steps when curating UnlearnCanvas and its key features.
Advantages of UnlearnCanvas. Next, we will highlight the advantages (A1-A3) of UnlearnCanvas to address the MU evaluation challenges (C1-C3) discussed in Sec. 3.
(A1) Style-object dual supervision enables a rich unlearning target bank for comprehensive evaluation. UnlearnCanvas offers a diverse range of unlearning targets by annotating each image with both style and object labels. This is in response to (C1). We have categorized the unlearning targets into two main domains: a style domain comprising $60$ artistic styles, and an object domain including $20$ object classes. This setup allows us to investigate the impact of style unlearning on the fidelity of generated objects in images and vice versa. Besides its role in MU evaluation, UnlearnCanvas could also serve as an annotated testbed for assessing style transfer accuracy and visual in-context learning performance; see details in Appx. C.
(A2) High stylistic consistency ensures precise style definitions and enables accurate quantitative evaluations. In UnlearnCanvas, images labeled with the same style exhibit high intra-style consistency and distinct inter-style differences, as exemplified in Fig. 2. This facilitates the distinction between different styles, effectively addressing challenge (C2). In a similar study to that presented in Tab. 2, we observe that the styles within UnlearnCanvas can be readily classified using ViT-Large (Dosovitskiy et al., 2020), achieving an accuracy of $98\%$ for images generated by SD finetuned on UnlearnCanvas. We refer readers to Tab. A2 for detailed results.
Figure 3: Evaluation metric illustration, with unlearning the ‘Van Gogh style’ as a case study. ✓ and ✗ denote desired and undesired results post unlearning.
(A3) Enabling in-depth retainability analyses for MU evaluation. Using UnlearnCanvas, we can assess DMs’ retainability post-unlearning from two perspectives: in-domain retainability, which evaluates the model’s performance on test-time generations within the same domain as the unlearning target, and cross-domain retainability, which assesses the model’s ability to generate concepts from a different domain. For example, as depicted in Fig. 3, we can define the ‘domain’ as either ‘artistic style’ or ‘object class’. If the unlearning target is the ‘Van Gogh’ style (i.e., the concept to be erased), then generating images in all styles except ‘Van Gogh’ but within the same object class (i.e., ‘Dog’ in Fig. 3) represents the in-domain case. Conversely, out-of-domain refers to generating images with different object classes (e.g., ‘Horses’ in Fig. 3), i.e., objects different from the one associated with the unlearning target.
Figure 4: Image examples with the label “Van Gogh Style” from WikiArt (Phillips & Mackintosh, 2011) and UnlearnCanvas. Images of the same artistic style in UnlearnCanvas exhibit high stylistic consistency compared to WikiArt.
Figure 5: An illustration of the evaluation pipeline proposed in this work using UnlearnCanvas when unlearning a specific target concept ‘Van Gogh Style’. Unlearning performances (including the unlearning effectiveness and retainability) are quantitatively assessed (marked in blue) to accurately reflect the unlearning performance portrait of each method. The unlearning target of the pipeline could traverse all the styles and objects to achieve a complete and comprehensive evaluation.
UnlearnCanvas vs. WikiArt.
To the best of our knowledge, WikiArt (Nichol, 2016) is the most relevant baseline dataset to ours. In Tab. A1, we provide a direct comparison of the key attributes of these two datasets. UnlearnCanvas differs from WikiArt in the following aspects. First, UnlearnCanvas includes a greater number of high-resolution images ($15$M) compared to WikiArt ($2$M), a factor that may enhance the training of state-of-the-art DMs. Second, UnlearnCanvas surpasses WikiArt in terms of both intra-style coherence and inter-style distinctiveness, as illustrated in Fig. 4, where images labeled with ‘Van Gogh Style’ from both datasets are compared. In UnlearnCanvas, the images exhibit high stylistic consistency, while WikiArt lacks the necessary clarity for precise assessment. This will hamper the MU evaluation as discussed in the challenges (C2) and (C3).
5 Evaluation Pipeline
In this section, we introduce the evaluation pipeline and the benchmarked MU methods for UnlearnCanvas. Fig. 5 illustrates the proposed evaluation pipeline for MU in DMs with a designated unlearning target. It comprises four phases (I-IV) to evaluate unlearning effectiveness, retainability, generation quality, and efficiency.
Phase I: Testbed preparation (Fig. 5a). Prior to unlearning, we commence by fine-tuning a specific DM, given by SD (StableDiffusion) v1.5, on UnlearnCanvas for text-to-image generation. We also train a vision transformer (ViT-Large) (Dosovitskiy et al., 2020) for the purposes of style and object classification after unlearning. We ensure that both the SD model and classifiers attain high performance in image generation and classification across all styles and objects. See detailed testbed preparation in Appx. A.
Phase II: Machine unlearning (Fig. 5b). Next, we utilize the selected MU methods for benchmarking to update the DM acquired in Phase I, aiming to eliminate a designated unlearning target (e.g., ‘Van Gogh style’). This results in the unlearned model (denoted as ${\bm{\theta}}_{\mathrm{u}}$), which is the DM updated to avoid generating the unlearning target. This unlearned DM will be used for unlearning performance evaluation.
Phase III: Answer set generation (Fig. 5c). Further, we utilize the unlearned model ${\bm{\theta}}_{\mathrm{u}}$ to generate a set of images conditioned on both the unlearning-related prompts and other innocent prompts. To comprehensively evaluate unlearning performance, three types of answer sets are generated:
Unlearning effectiveness: Images are generated in response to prompts that involve the unlearning target (e.g., any prompt like ‘An image in Van Gogh style’ for the case in Fig. 5). A successfully unlearned model will avoid generating images that match the unlearning target.
In-domain retainability: Images are generated with prompts in the same ‘domain’ as the unlearning target (e.g., style-related prompts for style unlearning), where the prompts can cover all the other styles provided in UnlearnCanvas except the one to be unlearned, e.g., ‘Van Gogh style’ in Fig. 5. This assesses the unlearned model’s ability to retain generation quality within the same domain.
Cross-domain retainability: Images are generated with prompts in other domains. For instance, in the case of style unlearning as shown in Fig. 5, the object domain is considered separate from the style domain. Prompts related to objects are used to evaluate the DM’s cross-domain generation capability after style unlearning.
Table 3: Performance overview of different MU methods evaluated with UnlearnCanvas dataset. The performance assessment includes unlearning accuracy (UA), in-domain retain accuracy (IRA), cross-domain retain accuracy (CRA), and FID. Symbols $\uparrow$ or $\downarrow$ denote whether larger or smaller values represent better performance. Results are averaged over all the style and object unlearning cases. The best performance regarding each metric is highlighted in bold.
Method Effectiveness Efficiency
Style Unlearning Object Unlearning FID ($\downarrow$) Time Memory Storage
UA ($\uparrow$) IRA ($\uparrow$) CRA ($\uparrow$) UA ($\uparrow$) IRA ($\uparrow$) CRA ($\uparrow$) (s) ($\downarrow$) (GB) ($\downarrow$) (GB) ($\downarrow$)
ESD (Gandikota et al., 2023a) $\mathbf{98.58}\%$ $80.97\%$ $93.96\%$ $92.15\%$ $55.78\%$ $44.23\%$ $65.55$ 6163 17.8 4.3
FMN (Zhang et al., 2023a) $88.48\%$ $56.77\%$ $46.60\%$ $45.64\%$ $90.63\%$ $73.46\%$ $131.37$ $\mathbf{350}$ 17.9 4.2
UCE (Gandikota et al., 2023b) $98.40\%$ $60.22\%$ $47.71\%$ $\mathbf{94.31}\%$ $39.35\%$ $34.67\%$ $182.01$ 434 $\mathbf{5.1}$ $\mathbf{1.7}$
CA (Kumari et al., 2023) $60.82\%$ $\mathbf{96.01}\%$ $92.70\%$ $46.67\%$ $90.11\%$ $81.97\%$ $\mathbf{54.21}$ 734 10.1 4.2
SalUn (Fan et al., 2023) $86.26\%$ $90.39\%$ $\mathbf{95.08}\%$ $86.91\%$ $\mathbf{96.35}\%$ $\mathbf{99.59}\%$ $61.05$ 667 30.8 4.0
Phase IV: MU performance assessment (Fig. 5d). After Phase III, the answer set undergoes style/object classification for unlearning performance assessment. This classification results in three quantitative metrics ①-③.
Unlearning accuracy (UA): It represents the proportion of images generated by the unlearned DM using the unlearning target-related prompt (e.g., the Van Gogh style prompt in Fig. 5), which are not correctly classified into the corresponding class for the case. Therefore, a higher UA indicates better unlearning performance in preventing image generation from the unlearning target-related prompts.
In-domain retain accuracy (IRA): It is given by the classification accuracy of images generated by the unlearned DM using innocent prompts (not relevant to the unlearning target) in the same domain, e.g., all styles other than the Van Gogh style in Fig. 5.
Cross-domain retain accuracy (CRA): Similar to IRA, it is given by the classification accuracy of images generated by the unlearned DM using innocent prompts but in different domains, e.g., all object categories in Fig. 5.
In addition to the above accuracy metrics, we also utilize the FID score (④) to evaluate the distribution-wise generation quality of the unlearned DM. Furthermore, we will also monitor the efficiency of the unlearning process, taking into account factors such as run-time (⑤), storage space requirements (⑥), and memory costs (⑦). In total, the performance evaluation involves 7 metrics (①-⑦).
MU methods for DMs to be benchmarked. In this work, we assess the 5 most recently proposed MU methods for DMs, including ESD (Gandikota et al., 2023a), FMN (Zhang et al., 2023a), UCE (Gandikota et al., 2023b), CA (Kumari et al., 2023), and Salun (Fan et al., 2023). We remark that SA (Heng & Soh, 2023), a recently proposed method, is excluded from our evaluation due to its excessive computational resource requirements and time consumption. Detailed training settings for each unlearning method can be found in Appx. A.
Figure 6: Performance visualization of Table 3. For UA, IRA, and CRA, the results are averaged over all the style and object unlearning scenarios. Other metrics undertake the inverse operation as a smaller values represent better performance. Results are normalized to $0\%\sim 100\%$ per metric.
Figure 7: Left: Heatmap visualization of the unlearning and retain accuracy of ESD on UnlearnCanvas. The $x$-axis shows the tested concepts for image generation using the unlearned model, while the $y$-axis indicates the unlearning target. Concept types are distinguished by color: styles in blue and objects in orange. The figure is separated into different regions to represent corresponding evaluation metrics and the unlearning scopes ($A$ for UA, $B$ for IRA, $C$ for CRA; ‘1’ for style unlearning, ‘2’ for object unlearning). Diagonal regions ($A1$ and $A2$) indicate unlearning accuracy, and off-diagonal values ($B1$, $B2$, $C1$, and $C2$) represent retain accuracy. Higher values in lighter color denote better performance. The first row serves as a reference for comparison before unlearning. Zooming into the figure is recommended for detailed observation. Right: Representative cases illustrating each region with images generated before and after unlearning a specific concept.
6 Experiment Results
In this section, extensive experiments are conducted on UnlearnCanvas following the evaluation pipeline in Sec. 5. We show that previous MU evaluation, primarily focused on unlearning effectiveness, may present a biased view without considering retainability assessment (Tab. 3). We also find that preserving cross-domain retainability, as quantified by CRA, is more challenging than in-domain retainability (measured by IRA), a previously overlooked aspect in the literature. Furthermore, we show that no single MU method consistently excels across all metrics, emphasizing the potential for advancements in existing methods (Fig. 6). In addition, we present a detailed case study on ESD, showcasing how our benchmark offers a comprehensive analysis of an unlearning method’s mechanism on forgetting and retaining information (Fig. 7). Last but not least, we show that different MU methods display distinct unlearning characteristics, which are inherently linked to their underlying unlearning mechanisms (Fig. 8).
Overall performance assessment of MU for DMs. In Tab. 3, we provide an overview of the performance of existing unlearning methods, using the metrics discussed in Sec. 5. The key results are summarized below.
(1) Retainability is essential for comprehensive assessment of MU in DMs. Previous evaluations heavily relied on UA (unlearning accuracy), but our assessment reveals a crucial insight: relying solely on UA can provide a skewed view of MU method performance. For example, when unlearning artistic painting styles (that we call style unlearning), ESD and UCE achieve similar UA (both around $98\%$), but their retainability (IRA and CRA) differs significantly, with one over $80\%$ and the other around $60\%$. Similar conclusions apply to unlearning an image object (that we call object unlearning). These findings emphasize the importance of including retainability metrics in the evaluation framework, as demonstrated by UnlearnCanvas.
(2) CRA is harder to retain than IRA. In both style and object unlearning, CRA presents a greater challenge for MU methods compared to IRA. This trend is consistent across various methods. For example, the UCE method (Gandikota et al., 2023b) exhibits a significant gap between its IRA and CRA, both in style ($60.22\%$ vs. $47.71\%$) and object ($39.35\%$ vs. $34.67\%$) unlearning scenarios. Similar patterns are observed with other methods like FMN and CA. This pattern suggests that while MU methods are somewhat effective at preserving concepts within the same domain (in terms of IRA), they face greater challenges in maintaining performance on unlearning target-unrelated concepts across domains (CRA), which has been overlooked in previous studies due to the absence of a multi-label dataset like UnlearnCanvas for systematic unlearning analysis.
(3) One MU method can perform differently in different domains. Take the FMN method (Gandikota et al., 2023b) as an example. It exhibits an evident performance difference between the two domains, achieving a UA of $88.48\%$ in style unlearning but a much lower UA of $45.64\%$ in object unlearning. In addition, its IRA and CRA metrics show a notable gap, with an IRA of $90.63\%$ for object unlearning compared to $56.77\%$ for style unlearning. These examples highlight that a method’s effectiveness in unlearning is not consistent across different scopes.
(4) No single method can excel in all aspects. Each method exhibits unique strengths and weaknesses. For example, while ESD (Gandikota et al., 2023a) typically leads an improvement in UA, it may lag behind in terms of IRA and CRA, e.g., in the case of object unlearning. Also, the SalUn method (Fan et al., 2023) shows exceptional retainability, with a CRA of $99.59\%$ in object unlearning, but its UA in the same category is lower at $86.91\%$. Furthermore, the UCE method (Kumari et al., 2023) enjoys remarkably high UA and efficiency but suffers from degradation in FID and retainability. For better visualization, in Fig. 6, we show this pattern by plotting the performance of different metrics reported in Tab. 1, illustrating that no method dominates across all radar chart dimensions. This underscores the complexity of unlearning in generative models and the need for further advancement.
Refer to caption Refer to caption Refer to caption
(a) ESD (b) SalUn (c) Illustration
Figure 8: Visualization of the unlearning directions of (a) ESD and (b) SalUn. This figure illustrates the conceptual shift of the generated images of an unlearned model conditioned on the unlearning target. Images generated by the post-unlearning models are classified and used to understand this shift. Edges leading from the object in the left column to the right signify that images generated conditioned on unlearning targets are instead classified as the shifted concepts after unlearning. This reveals the primary unlearning direction for each unlearning method. The most dominant unlearning direction for an object is visualized. Figure (c) provides visualizations of generated images using the prompt template ‘A painting of {object} in Sketch style.’ with object being each unlearning target.
A closer look into retainability: A case study on ESD. We next perform a more in-depth analysis of DMs’ retainability after unlearning, with a focus on the ESD method (Gandikota et al., 2023a), which has shown superiority in Tab. 3 for style unlearning. In Fig. 7 (left), we present a heatmap that illustrates the per-style/object unlearning accuracy (diagonal) and retain accuracy (off-diagonal) for ESD. The heatmap segments are labeled to indicate different regions of interest (ROIs) related to our evaluation metrics: A1 and A2 for Unlearning Accuracy (UA) in style and object unlearning, B1 and B2 for In-Domain Retain Accuracy (IRA) in style and object unlearning, and C1 and C2 for Cross-Domain Retain Accuracy (CRA) in style and object unlearning. We highlight the following observations.
First, we observe distinct behaviors of ESD in unlearning styles compared to objects. The ROI C2 appears darker than C1, suggesting a reduced ability to retain styles during the unlearning of objects. Similar patterns are observable between B2 and B1. These findings indicate that an MU method might demonstrate a preference within a specific domain but face challenges in others. For a clearer explanation, Fig. 7 (right) provides image generation examples of the DM before and after unlearning within different ROIs.
Second, we observe that style/object unlearning is relatively easier compared to retaining the generation performance of unlearned DMs conditioned on unlearning-unrelated prompts. This is evident in the contrast between ROIs A1/A2 and ROIs B1/B2 or C1/C2. One exception is the case of unlearning the object ‘Sea’, which exhibits relatively low UA in A2. This may be attributed to the image generation post-unlearning still resembling sea waves, as shown by the image examples in Fig. 7 (right) for A2.
Third, we present a similar accuracy heatmap for SalUn in Fig. A1. In comparison to ESD, SalUn demonstrates more consistent performance across different unlearning scenarios. However, SalUn does not achieve the same level of UA as ESD, as evidenced by the lighter diagonal values in Fig. A1. This observation underscores the trade-off between unlearning effectiveness and retainability of MU in DMs.
Understanding unlearning method’s behavior via unlearning directions. As noted earlier, different unlearning methods display distinct unlearning behaviors. To gain insights into the underlying reasons for these differences, Fig. 8 (a) and (b) visualize the ‘unlearning directions’ for ESD and SalUn, respectively. These unlearning directions are determined by connecting the unlearning target with the predicted label of the generated image from the unlearned DM conditioned on the unlearning target. As shown in Fig. 8 (a), ESD demonstrates a focused shift in image generation after object unlearning, with a predominant transition towards generating images labeled by ‘Sea’ and ‘Trees’. This behavior arises from ESD’s optimization process, designed to steer the generation of the DM away from a predefined concept. Consequently, images generated by the ESD-induced unlearned model consistently lack clearly identifiable objects, resembling waves and trees, which leads to their classification into the ‘Sea’ and ‘Trees’ classes; see Fig. 8 (c) for examples of generated images. In contrast, SalUn exhibits a more diverse range of unlearning directions, shifting images to 11 different objects. This diversity results from SalUn’s requirement to replace the unlearning target with a random concept. As shown in Fig. 8 (c), images generated by SalUn post-object unlearning still maintain some object contours (different from the original unlearning target) and better retain style information compared to ESD.
Expanding applications to style transfer. In addition to evaluating DMs’ unlearning performance, it’s worth noting that UnlearnCanvas has the potential to extend its utility to another generative modeling task: style transfer (Liu et al., 2021; Huang & Belongie, 2017; Li et al., 2019; Chen & Schmidt, 2016; Park & Lee, 2019; An et al., 2021). One significant challenge in evaluating style transfer methods is the absence of ground truth images for the given reference style image and content image (Cai et al., 2023). Thus, quantitative evaluations are also very limited in the literature. In Appx. C, we utilize UnlearnCanvas to assess 9 latest style transfer methods, providing enriched and precise evaluations. While not the primary focus of this work, this illustrates the great potential of UnlearnCanvas to contribute broader visual generative modeling tasks.
7 Conclusion
In this paper, we systematically reviewed the existing the stateful MU (machine unlearning) methods on DMs (diffusion models), and pinpointed the key challenges rooted in their evaluation systems, which could lead to incomplete, inaccurate, and biased assessments. In response, we propose UnlearnCanvas, a high-resolution stylized image dataset facilitating to build a comprehensive evaluation system for MU methods, together with a novel systematic retainability measurements, that have been long overlooked but are essential in a all-round evaluation. By benchmarking 5 stateful MU methods, we reveal novel insights into the their strengths and weaknesses, deepen the understanding on the underlying mechanism, and further illuminate several promising research directions. In the end, with a case study on style transfer, we show UnlearnCanvas enjoys high versatility and great potential to benchmark other generative modeling tasks beyond MU.
Impact Statement
This work helps improve the assessment and further promotes the advancement of MU (machine unlearning) methods for DMs (diffusion models), which are known to be effective in relieving or mitigating the various negative societal influences brought by the prevalent usage of DMs, which include but are not limited to the following aspects.
$\bullet$ Avoiding Copyright Issues. There is an urgent need for the generative model providers to scrub the influence of certain data on an already-trained model. In January 2023, a notable lawsuit targeted two leading AI art generators, Stable Diffusion (Rombach et al., 2022) and Midjourney (Midjourney et al.,, 2023), for alleged copyright infringement. Concurrently, incidents with the recently released Midjourney V6 (Midjourney et al.,, 2023) also highlighted a visual plagiarism issue on famous film scenes. These instances illuminate the broad copyright challenges inherent in the way of training data collection method of those foundation generative models’ training datasets. MU methods can be used as an effective method to remove the influence of the private data and avoid unnecessary retraining.
$\bullet$ Mitigating biases and stereotypes. Generative AI systems are known to have tendencies towards bias, stereotypes, and reductionism, when it comes to gender, race and national identities (Turk, 2023). For example, a recent study on the images generated with Midjourney revealed, that images associated with higher-paying job titles featured people with lighter skin tones, and that results for most professional roles were male-dominated (Vinker et al., 2023). MU is known to be effective in eliminating biases rooted in the training data. Moreover, UnlearnCanvas offers a flexible framework to benchmark MU techniques against bias removal, allowing for the creation and quantitative control of biases across different object classes for comprehensive bias removal studies.
Appendix
A Detailed Experiment Settings
A.1 Fine-tuning Style and Object Classifier with UnlearnCanvas.
Style and object classifiers need to be trained as part of the test bed proposed in our evaluation pipeline (Fig. 5). Here, we adopted a pretrained ViT-L/16 model (Dosovitskiy et al., 2020) pretrained on ImageNet and fine-tune it on UnlearnCanvas. UnlearnCanvas are split into the train set and test set with a ratio of $9:1$. After hyper-parameter tuning, the classifiers are trained with Adam optimizer with a learning rate of $0.01$ for 10 epochs. We noticed the model training accuracy can readily converge to $100\%$ even with 10 epochs of training.
A.2 Fine-tuning Stable Diffusion Model with UnlearnCanvas.
The other part of the testbed is a diffusion model capable of generating high quality images in all the styles associated with all the objects encompassed in UnlearnCanvas in order to guarantee a trustworthy and unbiased evaluation.
Training settings.
Practically, we fine-tune the pretrained StableDiffusion v1.5 on UnlearnCanvas for 20k steps with a learning rate of $61e-$ 6. Unless otherwise stated, we strictly follow the training configurations used in Stable Diffusion (Rombach et al., 2022). For each image in UnlearnCanvas, we annotate the data with text prompt An image of {object} in {style} style, where the object and style are the corresponding object and style label. For seed images, the style label we use is ‘photo’ style. We use the training scripts provided by Diffuser official tutorial (StableDiffusion) and the pretraining model card is runwayml/stable-diffusion-v1-5. During training, the checkpoints will be saved every 1000 steps.
Evaluation.
To evaluate the quality of the saved checkpoints and selcect the best one for unlearning study, the checkpoints are first used to generate an image set with the same prompt as training (An image of {object} in {style} style) by traversing all the possible style and object labels. Each prompt are used to generate $5$ images with different random seed. The image set for each checkpoint are fed into the style and object classifier trained in Appx. A.1. The model with the highest average performance on all the styles and objects are selected as the testbed for MU study. The classification performance are also used as a reference for later IRA/CRA comparison, which are disclosed in the first row of Fig. 7 (left). In Sec. 6, the experiments were built on the $50$ out of the $60$ styles, which served as the first edition of the dataset, and these styles are listed in Fig. 7. We remark that similar experiments can be done on all the 60 styles.
A.3 Implementation of the Tested Machine Unlearning Methods
In this work, we inspected 5 stateful MU methods for DMs. For each method, we use their publicly released source codes as code bases, which are listed below:
In particular, we adopt the following training settings to adapt the methods to our dataset:
  • ESD: based on the suggestions from the authors, ESD is used to only fine-tune the cross attention-related model weights (ESD-x). Other settings strict follow the ones used in the paper.
  • CA: in order to ablate concepts using CA, we first use ChatGPT to generate a list of simple prompts for each concept (including the styles and the objects), namely anchor prompts. Each anchor prompt is a simple one-sentence description of the unlearning target.
  • UCE: This method requires a guided concept (prompt) for each unlearning concept. For style unlearning, we use the prompt An image in {style*} style as the guided concept, where style* represents the next style in UnlearnCanvas in alphabetical order. Similarly, An image of object* is used for object unlearning, where object* is the next object in UnlearnCanvas in alphabetical order.
  • FMN: This method requires the images associated with the unlearning target. For simplicity and best performance, we randomly select 20 images associated with the unlearning concept. For the first stage of FMN, we run text inversion for 500 steps with a learning rate of $41e-$ 4, and for the second step, we used the inversed text to unlearn the cross attention layers of the model for $100$ steps. The hyper-parameters of learning rate, maximum steps, and tunable parameters (cross-attention or non-cross-attention) are carefully tuned with grid search.
  • SalUn: This method involves two steps, the mask finding (weight saliancy analysis) and the model unlearning. For mask finding, we tuned the mask ratio, while for unlearning, we tune the hyper-parameter learning rate and unlearning intensity. All the parameters are tuned with grid search. For both steps, the mask or model is trained with $10$ epochs.
B Additional Machine Unlearning Experiment Results
Table A1: Comparison with WikiArt, the most relevant dataset containing stylized images to ours. UnlearnCanvas stands out notably from WikiArt due to its characteristics of being supervised, balanced, and maintaining high stylistic cohesiveness.
Dataset
Resolution
(Pixels/Image)
Style-wise
Supervised
High Stylistic
Consistency
Class-wise
Balanced
WikiArt (Nichol, 2016) $\sim 2M$
Style-wise
Object-wise
UnlearnCanvas $\sim 15M$
Style-wise
Object-wise
Comparison to the most relevant dataset WikiArt.
We first make a comprehensive comparison between UnlearnCanvas and WikiArt on their key properties in Tab. A1. As we can see, UnlearnCanvas out performs WikiArt from the several perspectives. First, UnlearnCanvas contain images of much higher resolution (15M vs. 2M). This in particular benefits the training associated with DMs, where larger images in general help the model captures more details, such as the stylistic strokes. Second, the images in UnlearnCanvas are style-wise supervised. This means, for each seed image, a stylized counterpart can be find in each style class. This is beneficial for other tasks than MU, such as style transfer, which can provide a ground truth image for precise and robust evaluation. This will be detailed in Appx. C. Third, UnlearnCanvas enjoys high intra-style stylistic consistency and inter-style stylistic distinctiveness. This means the images of the same label share much more stylistic similarity, which is easier for a deep learning model to learn and to classify. This benefits can also be reflected by the training performance Appx. A.1 using UnlearnCanvas and WikiArt. The results are reported in Tab. 2 and Tab. A2, respectively. As we can see, the classifier is much more easily trained on UnlearnCanvas, justifying the higher discernible features within each style in UnlearnCanvas.
Table A2: Style classification results of a ViT-Large (Dosovitskiy et al., 2020) as a style classifier trained on UnlearnCanvas. After convergence, the classifier is tested on the test set and the image set generated by SD v1.5 finetuned on UnlearnCanvas.
UnlearnCanvas
Train Set
UnlearnCanvas
Test Set
Images by SD v1.5
tuned on UnlearnCanvas
Accuracy $100.0\%$ $99.9\%$ $98.8\%$
A fine-grained comparison per unlearning target: ESD vs. Salun.
Following the analysis of ESD, we next turn our focus to a comparative analysis with Salun, a method that demonstrated a better balance between unlearning and retaining according to Table 3. A similar accuracy heatmap for Salun is presented in Figure A1. Compared to ESD, Salun exhibits more consistent performance across various unlearning scenarios, as indicated by the more uniform color distribution in the heatmap. This also suggests enhanced retainability. However, it is noticeable that Salun does not reach the same level of UA (Unlearning Accuracy) as ESD, as evidenced by the darker diagonal values in Figure A1. This observation reinforces the existence of a trade-off between unlearning effectiveness and retainability in the visual generative MU task, a phenomenon paralleled in other tasks such as classification.
Figure A1: Heatmap visualization of SalUn. The plot setting is identical to Figure 7.
C Broader Use Cases of UnlearnCanvas
Although UnlearnCanvas is originally designed for benchmarking MU methods, we would like to demonstrate its broader use cases of benchmarking more generative modeling tasks, thanks to its good properties discussed in Sec. 4. In this section, we start with a case study on the task of style transfer, which is a much more well-studied topic than MU, but surprisingly also faces great challenges in building up a comprehensive and precise evaluation framework. In the next, we will first dissect the key challenges of the current style transfer evaluation framework, and then demonstrate how UnlearnCanvas efficiently resolves these challenges and further proposes a comprehensive and automated benchmark. Through extensive experiments, we draw demonstrate new insights from these results and illuminate the challenges of the future research directions. In the end, we will discuss the possibility of using UnlearnCanvas to benchmark more generative modeling tasks.
C.1 Benchmarking Style Transfer using UnlearnCanvas
Figure A2: An illustration of the task of style transfer.
Style transfer.
Style transfer is a long-standing topic and focuses on transferring the artistic style from one style image (also known as the reference image) $\mathbf{x}_{s}$ to a target content image $\mathbf{x}_{c}$. The most recent methods typically employ a neural network, denoted by ${\bm{\theta}}_{s}$, to extract style features and perform stylization in a single inference step, expressed as $\hat{\mathbf{x}}_{\mathrm{o}}=f_{{\bm{\theta}}_{s}}(\mathbf{x}_{s},\mathbf{x}_{c})$. Current state-of-the-art (SOTA) style transfer techniques exhibit remarkable generalization capabilities, successfully transferring styles not encountered during training and not requiring any further back-propagations. Figure 1 illustrates the pipeline of this task. Most existing literature (Liu et al., 2021; Huang & Belongie, 2017; Li et al., 2019; Chen & Schmidt, 2016; Park & Lee, 2019; An et al., 2021) utilize WikiArt (Phillips & Mackintosh, 2011) to provide different styles for training the stylization network ${\bm{\theta}}_{s}$. During the evaluation, the validation set of WikiArt will serve as the style (reference) images, together with the content images from the COCO dataset (Lin et al., 2014), to form a test bed for style transfer and style learning methods. However, such a evaluation scheme has some inherent limitations, which will be detailed below.
Issues and challenges in the evaluating methods for style transfer
Although the task of style transfer has been widely studied, its evaluations are still based on very limited quantitative or even sorely qualitative assessments, potentially leading to incomplete and inaccurate assessments (Cai et al., 2023). Upon examining the evaluation pipelines of over $10$ state-of-the-art (SOTA) style transfer methods, three significant challenges are identified within the current widely accepted evaluation frameworks. $\bullet$ Challenge I (C1): The lack of the ground truth images for style similarity evaluation. Unlike other vision tasks like classification (Singh & Singh, 2020), detection (Zou et al., 2023), and segmentation (Minaee et al., 2021), one of the key shortcomings of the current style transfer evaluation lies in the lack of the ground truth images $\mathbf{x}_{g}$ for the given reference style image $\mathbf{x}_{s}$ and the content image $\mathbf{x}_{c}$. Consequently, existing evaluation metrics, such as the style loss $\ell_{\mathrm{style}}$ (Gatys et al., 2015a, b), has to be calculated with the reference style image $\mathbf{x}_{s}$ as the ground truth $\mathbf{x}_{g}$, namely $\ell_{\mathrm{style}}(\mathbf{x}_{s},\hat{\mathbf{x}}_{o})$, rather than directly using the ground truth $\ell_{\mathrm{style}}(\mathbf{x}_{o},{\mathbf{x}}_{g})$. Obviously, such a indirect evaluation may lead to inaccurate results due to the different contents held in $\mathbf{x}_{s}$ and $\mathbf{x}_{o}$. Existing work has demonstrated that such evaluation metrics can lead to very different conclusion from that of the user study (An et al., 2021). Therefore, the creation of a supervised dataset with ground truth stylized images for every content image under each style is a timely remedy. $\bullet$ Challenge II (C2): The lack of algorithm stability evaluation against varied reference images $\mathbf{x}_{s}$. The evaluation of algorithm stability in style transfer has often been neglected. This assessment requires consistent performance of a method on a content image $\mathbf{x}_{c}$ across different style reference images ${\mathbf{x}_{s}}$ representing the same target artistic style. Ideally, an algorithm should maintain uniform quality across various references within a style, avoiding significant performance variations. Current challenges in such evaluations stem from the lack of reference sets that exhibit high stylistic consistency within each style. Figure 4 showcases examples from the widely-used WikiArt dataset. Despite sharing the same artistic label, images within a row show considerable divergence in visual appearance and style. Consequently, using these images as style references can lead to stable algorithms producing stylistically varied outputs, leading to misleading assessment results. Therefore, creating a dataset with high stylistic uniformity within each style category and clear differentiation between styles is crucial for accurate measurement of algorithm stability.
Refer to caption Refer to caption Refer to caption (a) Style Authenticity (b) Content Integrity (c) Algorithm Stability
Figure A3: An illustration of comprehensive performance metrics for style transfer tasks. (a) Style authenticity measures the style similarity between the generated image and the style image. (b) Content integrity measures the content preservation between the generated image and the content image. (c) Algorithm stability reflects the sensitivity of the algorithm to different style images. For all the metrics in the illustration, “Method II” is always better than “Method I”.
Building up a comprehensive evaluation pipeline with UnlearnCanvas for style transfer.
To revolutionize the evaluation framework with UnlearnCanvas, it is crucial to first understand the components that form a comprehensive evaluation pipeline, ensuring a thorough and impartial assessment of performance. In the realm of style transfer, three pivotal metrics stand paramount, as depicted in Figure,A3:
❶ Style Authenticity: Assesses the extent to which the style of the generated image aligns with that of the provided reference style image(s). ❷ Content Integrity: Measures how well the content features are preserved post style transfer. ❸ Algorithm Stability: Evaluates the algorithm’s robustness against variations in content subjects, target styles, or style reference image selections, while keeping other parameters constant.
In addressing the challenges (C1-C2), UnlearnCanvas proves to be inherently advantageous. First, the style-specific supervision embedded in UnlearnCanvas enables the provision of ground truth for quantitative evaluations of stylized images. Second, the extensive image collection within the same style category in UnlearnCanvas allows for the assessment of algorithm stability by applying style transfer methods to varied reference images. To encompass the aforementioned aspects of style transfer performance, we propose the following quantitative evaluation metrics:
① Style Loss: Utilizes a feature map-based style loss (Deng et al., 2022) to quantify the stylistic dissimilarity between image pairs, effectively representing the inverse of style authenticity. ② Content Loss: Employs a VGG-based, feature-map content loss (Deng et al., 2022; Zhang et al., 2022) to measure the visual dissimilarity between the reference and generated images, essentially mirroring content integrity. ③ Averaged Standard Deviation (STD): Computes the average STD of style and content loss w.r.t. the same reference image, reflecting algorithm stability.
Furthermore, similar to the MU task, we also consider efficiency metrics for each method, including:
④ Average Time Consumption: Measures the time required for performing style transfer. ⑤ Peak GPU Memory Consumption: Records the maximum GPU memory usage during the style transfer process. ⑥ Model Storage Memory Consumption: Assesses the memory requirement for storing the style transfer model.
With the introduction of evaluation metrics (①-⑥), we establish a comprehensive evaluation pipeline for style transfer. The process is as follows:
For each style transfer method, style transfer is executed within each object class. Specifically, for every style, images indexed from $1$ to $18$ serve as style reference images, while seed images corresponding to indices $19$ and $20$ are used as content images. Style and content loss are computed for each pair of style reference and content images, with the stylized images derived from the used seed content images serving as the ground truth for style loss. This results in a total of $60\times 20\times 18\times 2$ experimental trials for each method. For a specific style and content image pair, $18$ experiments are performed to calculate the Standard Deviation (STD) values for both style and content loss. These STD values are then averaged over all content images (amounting to $60\times 20\times 2$ cases). The findings from these comprehensive evaluations are presented in Table A3.
Following this evaluation pipeline, we scrutinized $9$ prominent style transfer methods, including SaNet (Park & Lee, 2019), Mcc (Deng et al., 2021), Mast (Huo et al., 2021), ArtFlow with its two variants (AF-AdaIn and AF-WCT) (An et al., 2021), IE-Contrast (Chen et al., 2021), Cast (Zhang et al., 2022), StyTr2 (Deng et al., 2022), and BLIP (Li et al., 2023).
Table A3: Performance overview of different style transfer methods evaluated with UnlearnCanvas dataset. The performance are assessed from the perspectives of stylistic authenticity (style loss), content integrity (content loss), algorithm stability (standard deviations from different dimensions), and efficiency. For all the metrics, smaller values are always preferred for better performance. The best performance per each metric is highlighted in bold. The standard deviations are first calculated with respect to different styles, object classes, or tested content images and then averaged in order to depict the algorithm stability from different perspectives.
Method Style Loss Content Loss Efficiency
Mean STD (Averaged over) Mean STD (Averaged over) Time (s/image) Memory (GB) Storage (GB)
Style Object Content Image Style Object Content Image
SaNet 23.48 2.73 2.87 1.87 0.85 0.12 0.17 0.09 0.29 2.3 0.11
Mcc 17.92 4.59 4.82 2.14 0.96 0.14 0.21 0.07 0.38 5.4 0.10
Mast 24.10 2.87 3.16 1.74 1.42 0.33 0.34 0.18 2.86 4.8 0.16
AF-AdaIn 20.78 2.96 3.13 1.65 1.09 0.11 0.19 0.05 0.53 6.3 0.08
AF-WCT 20.22 2.94 3.19 1.75 1.02 0.12 0.18 0.05 0.53 6.3 0.08
IE-Contrast 21.27 3.01 3.32 2.05 1.08 0.29 0.31 0.15 0.05 3.8 0.11
Cast 24.01 2.78 2.90 1.35 1.38 0.32 0.40 0.16 0.32 6.7 0.19
StyTr2 19.75 3.04 3.30 1.91 0.62 0.10 0.12 0.04 0.58 3.9 0.21
BLIP 25.43 2.90 3.06 2.03 1.61 0.30 0.34 0.16 8.87 7.2 7.23
Experiment results analysis.
Tab. A3 provides a systematic evaluation of the performance of various methods tested. From the analysis, we can derive several crucial insights:
First, it is evident that no single method excels across all evaluation metrics. Notably, Mcc demonstrates superior performance in maintaining stylistic authenticity, as indicated by a low style loss. Conversely, StyTr2 stands out in preserving content integrity, reflected by its minimal content loss.
Second, the assessment of standard deviation is indispensable for a comprehensive evaluation. The method with the optimal performance does not necessarily exhibit the greatest stability. This is particularly apparent in the style loss evaluation, where Mcc, despite achieving the best result in terms of style loss, exhibits the least stability, denoted by the highest standard deviation.
C.2 Broader Use Cases of UnlearnCanvas
In the preceding section, we demonstrated the application of UnlearnCanvas in refining evaluation metrics and frameworks for style transfer. Beyond this, we recognize the potential of UnlearnCanvas in diverse domains. Here, we delve into two illustrative examples:
Bias mitigation. Bias mitigation in DMs, which are now gaining popularity, can also benefit from UnlearnCanvas. Its hierarchical and balanced architecture enables the deliberate introduction of artificial biases by selectively omitting data from specific groups. For instance, by predominantly excluding images from styles other than the ‘Van Gogh Style’ within the ‘Dogs’ class, DMs fine-tuned on this dataset will inherently exhibit a tendency to generate images of dogs in the Van Gogh style, particularly when the style is not explicitly specified in the prompt. This approach not only allows for the manipulation and quantification of biases but also paves the way for UnlearnCanvas to become a standardized benchmark for bias mitigation, similar to the role of MU for DMs.
Vision in-context learning (V-ICL). V-ICL (Wang et al., 2023b, a; Bai et al., 2023; Wang et al., 2024) is another domain where UnlearnCanvas can be effectively applied. The field of V-ICL is in urgent need of robust, comprehensive methods for the fair assessment of existing models. In this context, the image pairs from UnlearnCanvas are ideally suited for evaluating various tasks such as style transfer, image inpainting, and image segmentation, offering a rich resource for nuanced and quantitative analyses.